# Dialogue Language Model with Large-Scale Persona Data Engineering

### v1

@@ Abstract
Mengze Hong1, Chen Zhang1, Chaotao Chen2, Rongzhong Lian2, Di Jiang2 1Hong Kong Polytechnic University 2WeBank Co. Ltd.
@@ 1 Introduction
The open-domain dialogue systems have gained significant interest due to their wide industrial applications, such as intelligent customer service agents, virtual assistance, and social chatbots Mandamadiotis et al. (2021); Zhou et al. (2024); Hong et al. (2024); Song et al. (2021b). Inspired by the recent success of large-scale pre-training in natural language processing, many neural dialogue models resort to pre-training on large-scale dialogue datasets Xu and Zhao (2021) and demonstrate substantial progress in open-domain dialogue. Notable examples include DialoGPT Zhang et al. (2020b), SPACE He et al. (2022), and Blender Shuster et al. (2022) for English dialogue, as well as CDial-GPT Wang et al. (2020), PLATO2 Bao et al. (2022), and EVA Gu et al. (2023) for Chinese.
While these methods can generate fluent and coherent responses, maintaining context consistency - particularly persona consistency - remains a common challenge that can lead to negative experiences in real-world human-bot interactions Welleck et al. (2019); An et al. (2024). As illustrated in Figure 1, the well-trained DialoGPT model struggles with maintaining persona consistency, i.e., revealing contradicted marriage status and generating out-of-character responses. Similar issues appear more frequently with the emerging LLM-based chatbots such as ChatGPT and Claud, where the response behavior often deviates from the instruction prompt and results in user dissatisfaction Dam et al. (2024); Song et al. (2024).
@@ 4.1 Model
The network architecture of PPDS is illustrated in Figure 3. Similar to the existing pre-training dialogue model, it employs Transformer blocks as the backbone. In order for efficient training on large-scale datasets, PPDS adopts the unified Transformer (also known as UniLM Dong et al. (2019)) instead of the typical encoder-decoder architecture for dialogue generation. By concatenating the persona, dialogue context, and response as a single input, the UniLM architecture can significantly reduce unnecessary computation of padding. Through the flexible mechanism of the self-attention mask, it can also simultaneously model the two tasks of dialogue context understanding and response generation with sharing parameters. Therefore, the UniLM architecture is more parameter-efficient than the encoder-decoder network Bao et al. (2021).
@@
In detail, the utterances in dialogue context are separated by a special token . The persona is represented by the concatenation of persona triple (i.e., ""), and each persona is also separated by a special token . Once we get the input representations, the UniLM Transformer will perform multi-head attention Vaswani et al. (2017) on the input to transform the embeddings into a sequence of hidden representations . Finally, we leverage Softmax to transform the hidden representations into the predictive probability of the target response Jiang et al. (2023). The details of the Transformer structure can refer to Dong et al. (2019).
@@ 4.3 Persona Augmentation
Since the large-scale persona dataset is constructed by extracting personas from responses, it may introduce a persona bias. Specifically, characters with extracted personas are often associated with persona-related responses. As a result, the model may generate persona-related responses whenever a persona profile is present, regardless of its relevance to the dialogue context.
To mitigate this bias, we propose augmenting the dialogue with unrelated personas, requiring the model to identify the relevant persona based on the dialogue context. Specifically, we collect all extracted personas and sample additional ones to supplement the persona profiles for each dialogue. If the sampled persona matches the existing persona type, we remove it to avoid potential contradictions. Finally, we merge this augmented dataset of unrelated personas with the raw persona dataset to eliminate the bias.
@@ 5.1 Experiment Setup
The following dialogue generation models are compared in the experiments.
Baseline Our vanilla PPDS, trained from scratch on PERSONA-CHAT without pre-training on the large-scale personal dialogue dataset.
DialoGPT is pre-trained on GPT-2 Radford et al. (2019) using Reddit comments. We compare its medium version, which reports the best performance Zhang et al. (2020b).
PPDS refers to our proposed model pre-trained on the large-scale persona dialogue dataset with persona augmentation.
PPDS-woP is pre-trained without a persona.
PPDS-woA is pre-trained on without persona augmentation.
DialoGPT-finetuned fine-tunes DialoGPT on PERSONA-CHAT by concatenating the persona with the dialogue context.
PPDS-finetuned is our PPDS pre-trained with persona augmentation and fine-tuned on PERSONA-CHAT.
PPDS-woP-finetuned is pre-trained without persona and fine-tuned on PERSONA-CHAT.
PPDS-woA-finetuned is pre-trained without persona augmentation and fine-tuned on PERSONA-CHAT.
@@ 5.2 Evaluation Metrics
We evaluate the response quality and persona consistency of the personal dialogue generation through both quantitative and human evaluations. For dialogue quality, we follow common practice Zhang et al. (2018) to employ the following quantitative metrics: (1) Perplexity (PPL). Lower perplexity means better language modeling. (2) Distinct 1/2 (Dist-1/2) Li et al. (2016a) denotes the ratio of distinct uni-grams/bi-grams, where higher distinct means better response diversity. (3) BertScore (BS) Zhang et al. (2020a) measures the coherence similarity between predicted response and target response measured through the BERT model. For persona consistency, we employ the ratios of responses that are entailed (E), neutral (N), and contradicted (C) to the personas, which are measured by a NLI model. We also calculate Consistency Score (CS) Madotto et al. (2019) to measure persona consistency, which summarizes the result of NLI as follows:
@@
As suggested in the empirical study Liu et al. (2016), the correlation between quantitative metrics and human judgments may be weak in open-domain dialogue generation. Therefore, we also conduct human evaluations in the experiments. Crowd-sourcing workers who are proficient in dialogue tasks are asked to evaluate the responses on the following metrics: (1) Fluency (Flu.) measuring whether the response is fluent and grammatically correct. (2) Coherence (Cohe.) measuring whether the response is relevant and consistent with the context. (3) Informativeness (Info.) evaluating whether the response is informative or not given the context. (4) Persona Consistency (P.C.) checking whether the response has conflicts with the persona, where means persona-related and consistent, means neutral, and means contradicted. Fluency, Coherence and Informativeness are all rated on a scale of [0, 1, 2], with higher scores indicating better quality.
5.3 Evaluation Result
6 Conclusion
Appendix A Qualitative Example
@@
| PPDS | Me too. My body would be so strong if I had a dry |
| wine! |

### v2

@@ Abstract
Mengze Hong1,2 Chen Jason Zhang1 Chaotao Chen2 Rongzhong Lian2 Di Jiang2††thanks: Di Jiang is the corresponding author: dijiang@webank.com 1Hong Kong Polytechnic University 2AI Group, WeBank Co., Ltd
@@ 1 Introduction
The open-domain dialogue systems have gained significant interest due to their wide industrial applications, such as customer service support Song et al. (2021b); Hong et al. (2024b), virtual assistance Mandamadiotis et al. (2021), and social chatbots Zhou et al. (2024); Ng et al. (2025). Inspired by the recent success of large-scale pre-training in natural language processing, many neural dialogue models resort to pre-training on large-scale dialogue datasets Xu and Zhao (2021) and demonstrate substantial progress in open-domain dialogue. Notable examples include DialoGPT Zhang et al. (2020b), SPACE He et al. (2022), and Blender Shuster et al. (2022) for English dialogue, as well as CDial-GPT Wang et al. (2020), PLATO-2 Bao et al. (2021), and EVA Gu et al. (2023) for Chinese.
While these methods can generate fluent and coherent responses, maintaining context consistency - particularly persona consistency - remains a common challenge that can lead to negative experiences in real-world human-bot interactions Welleck et al. (2019); An et al. (2024). As illustrated in Figure 1, the well-trained DialoGPT model struggles with maintaining persona consistency, i.e., revealing contradicted marriage status and generating out-of-character responses. Similar issues appear more frequently with the emerging LLM-based chatbots such as ChatGPT and Claude, where although the chatbot can comprehend user intention correctly Hong et al. (2024a), the response behavior often deviates from the instruction prompt and results in user dissatisfaction Tseng et al. (2024); Dam et al. (2024); Song et al. (2024).
@@ 4.1 Model
The network architecture of PPDS is illustrated in Figure 3. Similar to the existing pre-training dialogue model, it employs Transformer blocks as the backbone. In order for efficient training on large-scale datasets, PPDS adopts the unified Transformer (also known as UniLM Dong et al. (2019)) instead of the typical encoder-decoder architecture for dialogue generation. By concatenating the persona, dialogue context, and response as a single input, the UniLM architecture can significantly reduce unnecessary computation of padding. The flexible mechanism of the self-attention mask can also simultaneously model the two tasks of dialogue context understanding and response generation with sharing parameters. Therefore, the UniLM architecture is more parameter-efficient than the encoder-decoder network Bao et al. (2021). Additionally, UniLM has demonstrated promising performance across various downstream tasks Huang et al. (2022); Bao et al. (2020), highlighting its superiority and suitability.
@@
In detail, the utterances in dialogue context are separated by a special token . The persona is represented by the concatenation of persona triple (i.e., ""), and each persona is also separated by a special token . Once we get the input representations, the UniLM Transformer will perform multi-head attention Vaswani et al. (2017) on the input to transform the embeddings into a sequence of hidden representations . Finally, we leverage Softmax to transform the hidden representations into the predictive probability of the target response Jiang et al. (2023); Li et al. (2023). The details of the Transformer structure can refer to Dong et al. (2019).
@@ 4.3 Persona Augmentation
Since the large-scale persona dataset is built by extracting personas from the responses, there exists a persona bias in the raw persona dataset. Characters with extracted personas are often linked to persona-related responses, potentially misleading the model to generate such responses whenever a persona profile is present, regardless of its relevance to the dialogue context.
To mitigate this bias, we propose augmenting the dialogue with unrelated personas, compelling the model to identify the relevant persona based on the dialogue context. Specifically, we collect all extracted personas and sample some to supplement each dialogue’s persona profiles. If the sampled persona is of the same type as the existing one, we remove it to avoid contradictions. Finally, we merge this augmented dataset of unrelated personas with the raw persona dataset to eliminate the bias.
@@ 5.1 Experiment Setup
Baselines.
To evaluate the performance of PPDS, the following dialogue generation models including non-fine-tuned and fine-tuned ones are compared in the experiments.
Baseline: Our vanilla PPDS, trained from scratch on PERSONA-CHAT without pre-training on the large-scale personal dialogue dataset.
DialoGPT-finetuned: Fine-tuned DialoGPT on PERSONA-CHAT by concatenating the persona with the dialogue context.
PPDS: Our proposed model pre-trained on the large-scale persona dialogue dataset with persona augmentation.
PPDS-woP: Pre-trained without a persona.
PPDS-woA: Pre-trained without persona augmentation.
PPDS-finetuned: Our PPDS pre-trained on the large-scale persona dialogue dataset with persona augmentation and fine-tuned on the PERSONA-CHAT.
PPDS-woP-finetuned: PPDS fine-tuned on PERSONA-CHAT.
PPDS-woA-finetuned: Pre-trained without persona augmentation and fine-tuned on PERSONA-CHAT.
@@ 5.2 Evaluation Metrics
We evaluate the response quality and persona consistency of the personal dialogue generation through both quantitative and human evaluations. For dialogue quality, we follow common practice Zhang et al. (2018) to employ the following quantitative metrics: (1) Perplexity (PPL). Lower perplexity means better language modeling. (2) Distinct 1/2 (Dist-1/2) Li et al. (2016a) denotes the ratio of distinct uni-grams/bi-grams, where higher distinct means better response diversity. (3) BertScore (BS) Zhang et al. (2020a) measures the coherence similarity between predicted response and target response measured through the BERT model. For persona consistency, we employ the ratios of responses that are entailed (E), neutral (N), and contradicted (C) to the personas, which are measured by an NLI model. We also calculate Consistency Score (CS) Madotto et al. (2019) to measure persona consistency, which summarizes the result of NLI as follows:
@@
As suggested in the empirical study Liu et al. (2016), the correlation between quantitative metrics and human judgments may be weak in open-domain dialogue generation. Therefore, we also conduct human evaluations in the experiments. Crowd-sourcing workers who are proficient in dialogue tasks are asked to evaluate the responses on the following metrics: (1) Fluency (Flu.) measuring whether the response is fluent and grammatically correct. (2) Coherence (Cohe.) measuring whether the response is relevant and consistent with the context. (3) Informativeness (Info.) evaluating whether the response is informative or not given the context. (4) Persona Consistency (P.C.) checking whether the response has conflicts with the persona, where means persona-related and consistent, means neutral, and means contradicted. Fluency, Coherence, and Informativeness are all rated on a scale of [0, 1, 2], with higher scores indicating better quality.
@@
| PPDS | Me too. My body would be so strong if I had a dry wine! |
5.3 Evaluation Result
6 Conclusion
Beyond the discussed techniques, we also encourage exploration towards the following directions in constructing better persona datasets and training persona-consistent dialogue models for different application scenarios. First, the recent emergence of LLM-in-the-loop methodologies Hong et al. (2025) offers a promising path by incorporating the natural language understanding capabilities of LLMs to enhance the persona extraction process for complex, compositional personas. Second, the source of extraction can extend from textual data to multimodal datasets, particularly conversational speech that contains rich persona information Song et al. (2020c, 2022). Lastly, extending the persona from individual behaviors to larger entities, such as brand personality Aaker (1997), would further enhance the practical value of the proposed methods in various downstream domains, such as the hospitality and service sectors Ng et al. (2024).